OpenStudy (marieblossom16):

if I had 5-2 pies and some one ate 1 but I got 2 more how many pies do I have??

OpenStudy (alexandervonhumboldt2):

what does 5-2 mean? is it a subtraction?

OpenStudy (marieblossom16):

yep

OpenStudy (alexandervonhumboldt2):

ok first tell me 5-2 equal to ?

OpenStudy (marieblossom16):

3 :)

OpenStudy (alexandervonhumboldt2):

yeah how we have: \[(5-2)-1+2=3-1+2=?\]